### v1

@@ 2.1 Environment
We focus on the relationship between exporter $i$ and importer $j$ of an intermediate input. We denote by $\mathcal{Z}_{i}$ the set of importers connected to exporter $i,$ and by $\mathcal{Z}_{j}$ the set of exporters connected to importer $j$. These sets vary across firms and are treated as given.
Exporters and Supply
Exporter $i$ produces $q_{i}$ units of the unique input variety and sells them to all importers in $\mathcal{Z}_{i}$, where total output satisfies $q_{i}\equiv\sum_{j\in\mathcal{Z}_{i}}q_{ij}$, and $q_{ij}$ denotes the quantity of the intermediate input purchased by importer $j$.
We assume that exporter $i$ operates a short-run production technology with returns to scale parameter $\theta\in(0,1]$. Marginal cost (or equivalently, the short-run supply curve) is given by:
$c_{i}\equiv MC_{i}(q_{i})=k_{i}\,q_{i}^{\frac{1-\theta}{\theta}},$ (2.1)
where $k_{i}$ captures exogenous factors such as the exporter productivity or foreign wages.
This cost structure implies that the average cost of $q_{i}$ units is $\theta$-times the marginal cost. When $\theta<1$, indicating decreasing returns to scale, the short-run average cost lies below the marginal cost, and the gap between them widens as $q_{i}$ increases. In contrast, $\theta=1$ indicates a constant returns technology, with marginal and average costs constant and equal to $k_{i}$.
Importers and Demand
Importer $j$ combines domestic and foreign inputs, denoted by $q_{j}^{d}$ and $q_{j}^{f}$ respectively, to produce $q_{j}$ units of a final good, which is subsequently sold in a downstream market. The production technology of importer $j$ is given by:
$q_{j}=\varphi_{j}\left(q_{j}^{f}\right)^{\gamma}\left(q_{j}^{d\vphantom{f}}\right)^{\varrho-\gamma},$ (2.2)
where $\varphi_{j}$ is the importer’s productivity, $\gamma$ and $\varrho-\gamma$ are the output elasticities of foreign and domestic inputs, respectively, with $\varrho$ governing the degree of returns to scale in production. [^8] The domestic input $q_{j}^{d}$ can be interpreted as a constant returns to scale aggregator of primary factors, such as labor and domestic intermediates. Due to the lack of data on domestic input purchases, we model foreign and domestic intermediates as distinct rather than symmetric or substitutable.
The foreign input $q_{j}^{f}$ is a constant elasticity of substitution (CES) composite of differentiated input varieties sourced from exporters in the set $\mathcal{Z}_{j}$. Each variety $q_{ij}$ is weighted by a demand shifter $\varsigma_{ij}$, and the elasticity of substitution across varieties is given by $\rho>1$:
$q_{j}^{f}=\left(\sum_{i\in\mathcal{Z}_{j}}\varsigma_{ij}\,q_{ij}^{\frac{\rho-1}{\rho}}\right)^{\frac{\rho}{\rho-1}}.$ (2.3)
Downstream, each importer operates in a monopolistically competitive market, facing CES demand with constant elasticity $\nu>1$ and an importer-specific demand shifter $D_{j}$.
@@ 2.2 Bargaining Protocol and Gains From Trade
Each $i-j$ vertical interaction proceeds in two stages. In the first stage, the importer chooses input quantity $q_{ij}$ to minimize total input costs, taking the price $p_{ij}$ as given:
$\min_{q_{ij}}\;TC_{j}(q_{ij};p_{ij})=p_{ij}q_{ij}+\sum_{\ell\in\mathcal{Z}_{j}\setminus i}p_{\ell j}q_{\ell j}+p^{d}_{j}q^{d}_{j}.$
@@
The solution to this problem yields importer $j$’s input demand:
$q_{ij}=q_{j}^{f}\,\varsigma_{ij}^{\rho}\left(\frac{p_{ij}}{p_{j}^{f}}\right)^{-\rho}\quad\text{with}\quad q_{j}^{f}=\gamma\,c_{j}\,q_{j}\left(p_{j}^{f}\right)^{-1},$ (2.4)
where $p_{j}^{f}=\left(\sum_{i}\varsigma_{ij}^{\rho}p_{ij}^{1-\rho}\right)^{\frac{1}{1-\rho}}$ is the foreign input price index.
In the second stage, the importer and exporter determine the negotiated price $p_{ij}$ through Nash bargaining, taking the importer’s demand in equation (2.4) as given:
$\max_{p_{ij}}\left(\underbrace{\pi^{i}(p_{ij};q_{ij})-\tilde{\pi}^{i}_{(-j)}}_{GFT_{ij}^{i}}\right)^{1-\phi}\left(\underbrace{\pi^{j}(p_{ij};q_{ij})-\tilde{\pi}_{(-i)}^{j}}_{GFT_{ij}^{j}}\right)^{\phi}.$ (2.5)
In equation (2.5), $\phi\in(0,1)$ denotes the importer’s bargaining leverage, and $GFT_{ij}^{k}$ for $k\in\{i,j\}$ captures the gains from trade, defined as the difference between a firm’s profits with and without the bilateral relationship. In case of disagreement, each party falls back on trade with its other existing partners. All bilateral matches generate strictly positive surplus, and trade occurs in equilibrium across all relationships.
To solve equation (2.5), we adopt the Nash-in-Nash bargaining solution concept, in which each bilateral negotiation takes as given the outcomes of all other matches in the network (Horn and Wolinsky, 1988; Collard-Wexler et al., 2019). [^9] A common alternative is the sequential bargaining model of Stole and Zwiebel (1996), in which disagreement with one partner triggers renegotiation with others. This framework is suited to one-to-many settings such as labor markets, but is less appropriate for global supply chains, where firms engage in many-to-many relationships and isolated breakdowns rarely affect unrelated contracts. Moreover, it introduces additional complexity without yielding clear empirical gains in our context. For this reason, we leave the dependence on prices and quantities in other links in the network implicit throughout the analysis.
This setup implies two key properties. First, prices are allocative: the negotiated price $p_{ij}$ pins down the traded quantity $q_{ij}$. Second, the equilibrium allocation $(p_{ij},q_{ij})$ satisfies the importer’s demand in equation (2.4). We discuss the allocative implications of these assumptions, along with alternative bargaining protocols, in Section 2.5.
2.3 Equilibrium Prices and Allocations
The solution to (2.5) yields a bilateral price of the form $p_{ij}=\mu_{ij}\cdot c_{i},$ where $c_{i}$ is exporter $i$’s marginal cost, and $\mu_{ij}$ is a pair-specific markup.
We define two bilateral market shares that serve as key sufficient statistics in the model:
$s_{ij}$ $\equiv\frac{p_{ij}q_{ij}}{\sum_{k\in\mathcal{Z}_{j}}p_{kj}q_{kj}}\in[0,1],\quad\quad x_{ij}\equiv\frac{q_{ij}}{\sum_{k\in\mathcal{Z}_{i}}q_{ik}}\in[0,1],$
where $s_{ij}$ denotes exporter $i$’s supplier share, defined as its sales to importer $j$ as a fraction of $j$’s total foreign input expenditures, and $x_{ij}$ denotes importer $j$’s buyer share, equal to its purchases as a share of exporter $i$’s total output.
The following proposition characterizes the equilibrium markup as a convex combination of two limiting pricing regimes:
@@ 2.3 Equilibrium Prices and Allocations
The bilateral markup $\mu_{ij}$ can be expressed as
$\mu_{ij}=(1-\omega_{ij})\cdot\mu_{ij}^{\textnormal{oligopoly}}+\omega_{ij}\cdot\mu_{ij}^{\textnormal{oligopsony}},$ (2.6)
where the weight
$\omega_{ij}=\frac{\frac{\phi}{1-\phi}\lambda_{ij}}{1+\frac{\phi}{1-\phi}\lambda_{ij}}\in(0,1)$ (2.7)
represents the importer’s effective bargaining power, and $\lambda_{ij}\equiv-\frac{d\ln\pi_{j}}{d\ln(p_{ij}\,q_{ij})}\cdot\frac{\pi_{j}}{GFT_{ij}^{j}}$ captures the strength of the importer’s outside option.
@@
Equation (2.6) shows that the equilibrium markup lies between two polar cases. When the exporter holds all the bargaining power ($\phi\to 0$), the markup converges to the oligopoly case $\mu_{ij}^{\text{oligopoly}}$. When the importer holds full bargaining power ($\phi\to 1$), it converges to the oligopsony case $\mu_{ij}^{\text{oligopsony}}$. For intermediate values, the outcome reflects the influence of both parties, with weight $\omega_{ij}$ determined by bargaining power and the importer’s outside option (see Section 2.3.3).
We now examine these limiting cases in more detail and characterize how the bilateral markup depends on market shares and model primitives.
2.3.1 Oligopoly Markup
Under full exporter bargaining power ($\phi=0$), the markup takes the standard form:
$\mu_{ij}^{\text{oligopoly}}$ $=\frac{\varepsilon_{ij}}{\varepsilon_{ij}-1}\geq 1,$ (2.8)
where $\varepsilon_{ij}$ denotes the residual demand elasticity faced by exporter $i$, given by:
$\varepsilon_{ij}=(1-s_{ij})\cdot\rho+s_{ij}\cdot\eta.$ (2.9)
This elasticity is a weighted average of two components: the elasticity of substitution across foreign suppliers, $\rho$, and the elasticity $\eta$ of the importer’s foreign input bundle $q_{j}^{f}$ with respect to its price index $p_{j}^{f}$:
$\eta\equiv-\frac{d\ln q_{j}^{f}}{d\ln p_{j}^{f}}=\frac{(\varrho-\gamma)+\nu\cdot(1-(\varrho-\gamma))}{\varrho+\nu\cdot(1-\varrho)}.$
Provided that $\rho>\eta$, a standard parameter condition, the exporter’s markup (2.8) increases with its supplier share $s_{ij}$, reflecting greater oligopoly power.
This case corresponds to the firm-to-firm trade setting in Dhyne et al. (2022). It is also closely related to the oligopolistic competition model of Atkeson and Burstein (2008), but differs in three key respects. First, supplier shares are defined at the match level, rather than at the firm or industry level, reflecting our assumption of a fixed network. Second, whereas the outer nest elasticity $\eta$ is typically treated as a fixed preference parameter, here it summarizes how downstream market structure–determined by both technology ($\gamma$, $\varrho$) and demand ($\nu$)–shapes the bargaining environment. Third, we allow for decreasing returns to scale in production, in contrast to the constant returns assumption common in related models.
Figure 1 illustrates the equilibrium outcome for a representative $i-j$ match. The downward-sloping residual demand and marginal revenue curves (black and gray) intersect with the upward-sloping residual marginal and average cost curves (blue and red).
Panel (A) depicts the case with $\phi=0$. The equilibrium quantity is determined by the intersection of marginal cost and marginal revenue, and the price lies on the residual demand curve. Exporter rents have two components: the oligopoly rent (red), which corresponds to the markup in equation (2.8), and the quasi-rent (purple), which arises from the gap between marginal and average cost under decreasing returns to scale.
Figure 1: Illustration of Equilibrium Allocations With Different Bargaining Power
(a) $\phi=0$
(b) $\phi\in(0,1)$
(c) $\phi=1$
2.3.2 Oligopsony Markdown
Under full importer bargaining power ($\phi=1$), the bilateral markup simplifies to:
$\mu_{ij}^{\text{oligopsony}}=\theta\cdot\frac{1-(1-x_{ij})^{\frac{1}{\theta}}}{x_{ij}}\leq 1,$ (2.10)
which we refer to as an oligopsony markup, or markdown, since it lies weakly below one. [^10] Although $\mu_{ij}^{\text{oligopsony}}$ denotes a price–cost ratio, we refer to it as a markdown because it falls below one. This differs from standard oligopsony models, where markdowns typically reflect a wedge between input prices and marginal revenue product. See Section 2.5 for further discussion.
Expression (2.10) characterizes the lowest price at which the exporter is willing to supply, with the importer extracting the exporter’s entire surplus through bargaining. [^11] Formally, the markdown is proportional to the percentage increase in quasi-rents attributable to the $i-j$ match. See equation (A.9) in Appendix A.1. The markdown declines with the buyer’s share $x_{ij}$: larger buyers induce greater quasi-rents and can negotiate lower prices. As $x_{ij}\to 0$, the importer behaves atomistically, generates no quasi-rents, and pays a price equal to marginal cost.
This dependence of the price–cost ratio on $x_{ij}$ reflects a form of oligopsony power that arises only under decreasing returns to scale. When $\theta=1$, marginal and average costs coincide, quasi-rents vanish, and the importer can extract only oligopoly rents. In this case, the price cannot fall below marginal cost, and the markdown remains fixed at one.
Panel (C) of Figure 1 illustrates the equilibrium under $\phi=1$. Compared to Panel (A), the price lies at the intersection of the downstream demand curve and the exporter’s residual average cost curve, leaving no surplus to the exporter. The result is a lower price and higher quantity relative to the oligopoly case.
@@ 2.3.2 Oligopsony Markdown
The terms “buyer power” and “oligopsony power” are often used interchangeably in the literature, but rarely precisely defined (Noll, 2005). We distinguish between these concepts in our framework.
@@
We use buyer power or importer market power to refer broadly to the importer’s ability to influence prices through bargaining (i.e., any $\phi>0$). In contrast, we reserve oligopsony power for the specific case when a higher buyer share $x_{ij}$ enables the importer to negotiate lower prices, holding $\phi$ fixed. Importantly, oligopsony power requires both $x_{ij}>0$ and $\theta<1$, as it depends on the scale-dependent gap between marginal and average cost.
2.3.3 Two-Sided Market Power and Bargaining Weights
Panel (B) of Figure 1 shows the intermediate case with $\phi\in(0,1)$. The equilibrium price lies on the thick blue curve and reflects a weighted average of the outcomes in Panels (A) and (C), consistent with Proposition 1. The resulting price–marginal cost ratio can exceed or fall below one, depending on the relative bargaining power of the exporter and importer.
Determinants of $\omega_{ij}$.
The bargaining weight $\omega_{ij}$ in equation (2.6), defined in equation (2.7), governs the balance between oligopoly and oligopsony forces in price setting. It depends both on the exogenous bargaining power $\phi$ and an endogenous term $\lambda_{ij}$, defined as:
[…]
To interpret $\lambda_{ij}$, we decompose it into two terms. The cost exposure component, $\lambda_{ij}^{C}=\frac{(\eta-1)s_{ij}}{\varepsilon_{ij}-1}$, captures how sensitive importer $j$’s profits are to price changes from supplier $i$ and increases with the supplier share $s_{ij}$. The network dependence term, $\lambda_{ij}^{N}=\left(1-(1-s_{ij})^{\frac{\eta-1}{\rho-1}}\right)^{-1}$, reflects how much of importer $j$’s profits rely on the match with $i$, and declines with $s_{ij}$ as reduced diversification weakens the importer’s outside option.
Taken together, $\lambda_{ij}$ captures both the importer’s incentive to negotiate aggressively (through cost exposure) and the strength of its fallback option (through network dependence). It follows a hump-shaped pattern in $s_{ij}$, increasing at low values and declining at higher ones, and converges to 1 as $s_{ij}\to 0$ or $s_{ij}\to 1$.
This interpretation clarifies the meaning of $\omega_{ij}$ as a measure of the importer’s effective bargaining power. While $\phi$ governs baseline influence in Nash bargaining, $\lambda_{ij}$ endogenously adjusts it based on the structure of the trading relationship.
Testable Implications.
Equation (2.6) writes the bilateral markup $\mu_{ij}$ as a convex combination of the oligopoly markup and the oligopsony markdown, with bargaining power as weight. Since these components scale with bilateral market shares, the model delivers clear predictions on how markups co-move with $s_{ij}$ and $x_{ij}$ under two-sided market power.
Proposition 2.
The bilateral markup $\mu_{ij}$ exhibits the following properties:
  1. 1.
    Markup and Exporter’s Share. If $\mu_{ij}$ increases with $s_{ij}$, then $\phi<1$.
  2. 2.
    Markup and Importer’s Share. If $\mu_{ij}$ decreases with $x_{ij}$, then $\phi>0$ and $\theta<1$.
Proof: See Appendix A.2.
Proposition 2 provides sufficient conditions for detecting two-sided market power. A positive relationship between $\mu_{ij}$ and supplier share $s_{ij}$ implies $\phi<1$, since exporters retain pricing power and $\mu_{ij}^{\text{oligopoly}}$ increases with $s_{ij}$. If $\phi=1$, prices are fully disciplined by buyers, and no such pattern should emerge. While the dependence of $\omega_{ij}$ on $s_{ij}$ may attenuate this effect, a positive slope remains inconsistent with full buyer power.
Conversely, a negative relationship between $\mu_{ij}$ and buyer share $x_{ij}$ signals oligopsony power. This requires both positive bargaining leverage for importers ($\phi>0$) and decreasing returns to scale ($\theta<1$). When both conditions are met, markups reflect the joint influence of importer and exporter market power.
2.4 Equilibrium: Tariff Pass-Through Elasticities

### v2

@@ 2.1 Environment
We focus on the relationship between exporter $i$ and importer $j$ of an intermediate input. We denote by $\mathcal{Z}_{i}$ the set of importers connected to exporter $i$, and by $\mathcal{Z}_{j}$ the set of exporters connected to importer $j$. These sets vary across firms and are taken as given.
Exporters and Supply.
Exporter $i$ supplies a differentiated input to importers in $\mathcal{Z}_{i}$. Total exported quantity is $q_{i}=\sum_{j\in\mathcal{Z}_{i}}q_{ij}$, where $q_{ij}$ denotes importer $j$’s demand.
The exporter’s marginal cost is
$MC_{i}(q_{i})=k_{i}\,q_{i}^{\frac{1-\theta}{\theta}},$ (2.1)
where $\theta\in(0,1]$ governs returns to scale and $k_{i}$ captures exogenous cost shifters, such as productivity or foreign wages. When $\theta<1$, the technology exhibits decreasing returns, so marginal cost rises with output. When $\theta=1$, marginal cost is constant and equal to $k_{i}$. The implied total cost function is $TC_{i}(q_{i})=\theta\,k_{i}\,q_{i}^{\frac{1}{\theta}}$, so average cost equals $\theta$ times marginal cost.
Importers and Demand.
Importer $j$ produces a final good $q_{j}$ by combining a foreign input ($q_{j}^{f}$) with a domestic input ($q_{j}^{d}$) according to the production function
$q_{j}=\varphi_{j}\,(q_{j}^{f})^{\gamma}\,(q_{j}^{d})^{\varrho-\gamma},$ (2.2)
where $\varphi_{j}$ denotes productivity, $\gamma$ the output elasticity of foreign inputs, and $\varrho$ the returns-to-scale parameter. [^8] Since domestic input purchases are not observed, we impose a unit elasticity of substitution between foreign and domestic inputs, which fixes the cost share of $q_{j}^{d}$ and $q_{j}^{f}$. Under this restriction, $q_{j}^{d}$ can be interpreted as a constant-returns aggregator of primary factors such as labor or domestic intermediates.
The foreign input consists of a bundle of input varieties supplied by the exporters in $\mathcal{Z}_{j}$ and aggregated according to a CES technology:
$q_{j}^{f}=\left(\sum_{i\in\mathcal{Z}_{j}}\varsigma_{ij}\,q_{ij}^{\,\frac{\rho-1}{\rho}}\right)^{\!\frac{\rho}{\rho-1}},$ (2.3)
where $\varsigma_{ij}$ is a demand shifter, and $\rho>1$ is the substitution elasticity across input varieties.
Downstream, importer $j$ sells its final good in a monopolistically competitive market and faces CES demand with elasticity $\nu>1$. Revenue is therefore given by
$R_{j}(q_{j})=D_{j}\,q_{j}^{\frac{\nu-1}{\nu}},$ (2.4)
where $D_{j}$ denote an exogenous demand shifter.
@@ 2.2 Bargaining Protocol and Gains From Trade
Each bilateral interaction between importer $j$ and exporter $i$ unfolds in two stages. In the first stage, for a given output target $q_{j}$, importer $j$ chooses foreign inputs $\{q_{\ell j}\}_{\ell\in\mathcal{Z}_{j}}$ and domestic input $q_{j}^{d}$ to minimize total cost, taking input prices as given:
$\min_{\{q_{\ell j}\},\,q_{j}^{d}}\quad\sum_{\ell\in\mathcal{Z}_{j}}p_{\ell j}q_{\ell j}+p_{j}^{d}q_{j}^{d}\quad\text{s.t.}\quad\varphi_{j}(q_{j}^{f})^{\gamma}(q_{j}^{d})^{\rho-\gamma}\geq q_{j}.$ (2.5)
@@
Solving this problem yields the standard CES demand for input varieties:
$q_{ij}=\varsigma_{ij}^{\rho}\left(\frac{p_{ij}}{p_{j}^{f}}\right)^{-\rho}q_{j}^{f},$ (2.6)
where $q_{j}^{f}$ is the foreign input aggregator and $p_{j}^{f}$ the associated price index, satisfying $p_{j}^{f}q_{j}^{f}=\gamma\,c_{j}\,q_{j}$, with $c_{j}=MC_{j}(q_{j})$ denoting importer $j$’s marginal cost.
In the second stage, the two parties bargain over the price $p_{ij}$, taking the demand schedule (2.6) as given. [^9] This protocol is analogous to the employer “right-to-manage” model in labor economics (e.g., azkarate2025union), where wages are bargained while firms retain the choice of employment. Operating profits are given by $\pi^{i}_{ij}=\sum_{k\in\mathcal{Z}_{i}}p_{ik}q_{ik}-TC_{i}(q_{i})$ for the exporter and $\pi^{j}_{ij}=R_{j}(q_{j})-\sum_{\ell\in\mathcal{Z}_{j}}p_{\ell j}q_{\ell j}-p_{j}^{d}q_{j}^{d}$ for the importer, where $TC_{i}(q_{i})$ and $R_{j}(q_{j})$ are defined in Section 2.1. If bargaining fails and link $(i,j)$ is severed, each party falls back on profits from its remaining matches, denoted $\tilde{\pi}_{(-j)}^{i}$ and $\tilde{\pi}_{(-i)}^{j}$.
We define the gains from trade as $GFT_{ij}^{k}\equiv\pi^{k}_{ij}-\tilde{\pi}_{(-\cdot)}^{k}$ for $k\in\{i,j\}$, which measure the value each firm derives from the match. For the exporter, these gains correspond to revenue from $j$ net of the incremental cost of serving that buyer. For the importer, losing variety $i$ raises the input price index $p_{j}^{f}$, increasing marginal costs, reducing optimal output, and lowering downstream revenues. In our setting, these gains are strictly positive. [^10] Appendix B.3 discusses an extension with more general nonparametric outside options.
The negotiated price solves a Nash bargaining problem that splits these gains, with $\phi\in(0,1)$ denoting the importer’s exogenous bargaining weight:
$\max_{p_{ij},\,q_{ij}}\;\left(GFT_{ij}^{i}\right)^{1-\phi}\;\left(GFT_{ij}^{j}\right)^{\phi}\quad\text{s.t.}\quad q_{ij}=q_{j}^{f}\varsigma_{ij}^{\rho}\left(\frac{p_{ij}}{p_{j}^{f}}\right)^{-\rho}.$ (2.7)
We adopt the Nash-in-Nash solution concept, taking prices and quantities on all other links as given (Horn1988; collard2019nash).
2.3 Price Characterization
To characterize equilibrium prices, it is useful to summarize each bilateral relationship with two market shares that capture each party’s relative importance to the other. For a given link $(i,j)$, the exporter’s supplier share $s_{ij}$ is its share of the importer’s foreign input expenditure; the importer’s buyer share $x_{ij}$ measures its share of the exporter’s total output. Formally,
$s_{ij}\equiv\frac{p_{ij}q_{ij}}{\sum_{k\in\mathcal{Z}_{j}}p_{kj}q_{kj}},\qquad\text{and}\qquad x_{ij}\equiv\frac{q_{ij}}{\sum_{k\in\mathcal{Z}_{i}}q_{ik}}.$ (2.8)
We also define $\xi(GFT_{ij}^{j},p_{ij})$ as the elasticity of the importer’s gains from trade with respect to the bilateral price:
$\xi(GFT_{ij}^{j},p_{ij})\equiv-\frac{d\ln GFT_{ij}^{j}}{d\ln p_{ij}}=\frac{(\eta-1)s_{ij}}{1-(1-s_{ij})^{\frac{\eta-1}{\rho-1}}}.$ (2.9)
With these objects in hand, we can now present our first key theoretical result.
@@ 2.3 Price Characterization
The solution to (2.7) yields the price–cost ratio $\mu_{ij}\equiv p_{ij}/MC_{i}$, which satisfies
$\mu_{ij}=(1-\omega_{ij})\,\mu^{\textnormal{oligopoly}}(s_{ij})+\omega_{ij}\,\mu^{\textnormal{oligopsony}}(x_{ij}),$ (2.10)
with
$\omega_{ij}=\frac{\phi\,\xi(GFT_{ij}^{j},p_{ij})}{\phi\,\xi(GFT_{ij}^{j},p_{ij})+(1-\phi)(\varepsilon_{ij}-1)}\in(0,1).$ (2.11)
Proof. See Appendix A.1.
The first component in equation (2.10), $\mu^{\textnormal{oligopoly}}_{ij}$, is the standard oligopoly markup arising under Bertrand competition with differentiated products:
$\mu_{ij}^{\text{oligopoly}}=\frac{\varepsilon_{ij}}{\varepsilon_{ij}-1}\geq 1\qquad\text{where}\qquad\varepsilon_{ij}=(1-s_{ij})\rho+s_{ij}\eta.$ (2.12)
This expression coincides with the equilibrium price-cost ratio when the exporter has full bargaining power ($\phi=0$), in which case $\omega_{ij}=0$.
Here, the residual demand elasticity $\varepsilon_{ij}$ is a weighted average of the within-input elasticity $\rho$ and the elasticity
$\eta\equiv-\frac{d\ln q_{j}^{f}}{d\ln p_{j}^{f}}=\frac{(\varrho-\gamma)+\nu(1-(\varrho-\gamma))}{\varrho+\nu(1-\varrho)},$ (2.13)
which depends on technology ($\gamma,\varrho$) and demand ($\nu$). [^11] This is different than in standard nested-CES models with Bertrand competition, where $\eta$ is typically a primitive preference parameter (e.g., Atkeson2008). When $\rho>\eta$, the markup increases in the supplier share $s_{ij}$, reflecting greater oligopoly power.
The second component in equation (2.10), $\mu^{\textnormal{oligopsony}}_{ij}$, is given by
$\mu_{ij}^{\text{oligopsony}}=\theta\cdot\frac{1-(1-x_{ij})^{1/\theta}}{x_{ij}}\leq 1.$ (2.14)
This expression corresponds to the equilibrium price–cost ratio when the importer has full bargaining power ($\phi=1$), in which case $\omega_{ij}=1$. Since $\mu_{ij}^{\text{oligopsony}}\leq 1$, we refer to it as an oligopsony markdown. [^12] Although we refer to $\mu_{ij}^{\text{oligopsony}}$ as a markdown, it is formally a price–cost ratio. This differs from the conventional definition of a markdown as the wedge between the input price and the input’s marginal revenue product.
The oligopsony markdown has a natural quasi-rents interpretation. Under full importer bargaining power, the importer extracts all exporters’ surplus and bargain a price equals the exporter’s incremental cost of serving them. Expressed relative to current marginal cost, the corresponding price-cost ratio is
$\mu_{ij}^{\text{oligopsony}}=\frac{TC_{i}(q_{i})-TC_{i}(q_{i}-q_{ij})}{MC_{i}(q_{i})\,q_{ij}}.$
The numerator is the exporter’s actual cost saving if the relationship with buyer $j$ is severed. The denominator is what that cost would be if all units sold to $j$ were valued at current marginal cost. Under decreasing returns, the incremental cost of serving $j$ lies below current marginal cost, allowing the negotiated price to fall below marginal cost. This gap, which owes to quasi rents, captures oligopsony power in the model.
This markdown decreases with the buyer share $x_{ij}$. Larger buyers account for a greater share of the exporter’s output, so walking away from the relationship generates larger cost savings for the exporter and strengthens the importer’s bargaining position. As $x_{ij}\to 0$, the buyer is atomistic, the incremental cost of serving the match converges to marginal cost, and $\mu_{ij}^{\text{oligopsony}}\to 1$, so that $p_{ij}=MC_{i}$. As $x_{ij}\to 1$, the importer is a monopsonist, the negotiated price converges to average cost, and $\mu_{ij}^{\text{oligopsony}}\to AC_{i}/MC_{i}=\theta$. [^13] Our oligopsony mechanism is observationally equivalent to quantity discounts across a supplier’s buyers: conditional on a supplier and product, marginal cost is common across buyers, yet buyers that account for a larger share of the supplier’s output $(x_{ij})$ obtain higher markdowns and pay lower negotiated prices.
This mechanism operates only under decreasing returns to scale. When $\theta=1$, marginal and average costs coincide, leaving no surplus to extract. As a result, the negotiated price cannot fall below marginal cost, and the markdown in (2.14) equals one for all $x_{ij}$.
Bargaining Weights.
Equation (2.10) expresses the bilateral markup as a weighted average of the oligopoly markup (2.12) and the oligopsony markdown (2.14). The weight $\omega_{ij}$ governs the relative importance of these two forces. It is increasing in the exogenous bargaining parameter $\phi$, but it is also endogenous because it depends on both the importer’s gains-from-trade elasticity $\xi(GFT_{ij}^{j},p_{ij})$ and the residual demand elasticity $\varepsilon_{ij}$. Thus, even for a given value of $\phi$, effective bargaining power varies systematically across matches.
In particular, for $\phi\in(0,1)$, the weight $\omega_{ij}$ is hump-shaped in the supplier share $s_{ij}$. When $s_{ij}$ is small, the bilateral price has little effect on the importer’s overall costs, so the importer has little to gain from bargaining aggressively. When $s_{ij}$ is very large, the importer is highly exposed to the relationship but also depends heavily on it, which weakens its outside option. Effective bargaining power is therefore strongest at intermediate supplier shares, where cost exposure is significant, but substitution remains feasible.
Equilibrium Allocation.
Figure 1: Equilibrium Allocations With Different Bargaining Power
Figure 1 illustrates the equilibrium allocation for a single $i$-$j$ match under alternative bargaining weights $\phi$. The exporter faces a downward-sloping residual demand curve and its associated marginal revenue curve, shown in black and grey, respectively. The importer faces an upward-sloping residual marginal and average cost curves, shown in blue and red.
Panel (A) shows the oligopoly benchmark ($\phi=0$). Quantity is pinned down by the intersection of marginal revenue and marginal cost, and price is then determined from residual demand. Exporter rents can be decomposed into two components: oligopoly rents, shown in dark blue, equal to (pMC)q(p-MC)q, and quasi-rents, shown in light blue, equal to the gap between marginal and average cost under decreasing returns, (MCAC)q(MC-AC)q.
Panel (B) shows full importer bargaining power ($\phi=1$). The negotiated allocation lies at the intersection of residual demand and average cost curves, leaving the exporter with zero net gains from trade. In this case, the importer captures the full quasi-rents generated by decreasing returns. Relative to the oligopoly benchmark, the negotiated price is lower and the traded quantity is higher.
Panel (C) shows the intermediate case ($\phi\in(0,1)$). The negotiated outcome lies between the two polar regimes, consistent with Proposition 1. The striped area highlights the corresponding range of price–marginal cost ratios, which may lie either above or below one depending on bargaining power and the magnitude of quasi-rents. The grey shaded region represents total exporter surplus, equal to (pAC)q(p-AC)q, as bargaining power varies between the two extremes. As $\phi$ rises, the importer captures a larger share of the surplus, so exporter rents decline and converge to zero at $\phi=1$.
Markups and Market Shares.
Both components of the bilateral markup vary systematically with bilateral market shares. A higher supplier share $s_{ij}$ raises the oligopoly markup, while a higher buyer share $x_{ij}$ lowers the oligopsony markdown. The co-movements between prices and bilateral shares are thus informative about the presence of the two forces, as summarized in the following proposition.
Proposition 2.
The bilateral markup $\mu_{ij}$ satisfies:
  1. 1.
    Supplier Share. If $\mu_{ij}$ increases in $s_{ij}$, then $\phi<1$.
  2. 2.
    Buyer Share. If $\mu_{ij}$ decreases in $x_{ij}$, then $\phi>0$ and $\theta<1$.
@@
These cross-sectional predictions guide the empirical analysis in Section 3.4.
@@
We use buyer power or importer market power (equivalently, countervailing power) to refer broadly to the importer’s ability to influence prices through bargaining (i.e., any $\phi>0$). In contrast, we reserve oligopsony power for the case in which a higher buyer share $x_{ij}$ allows the importer to negotiate lower prices, other things equal. Oligopsony power therefore requires both $\phi>0$ and $\theta<1$, as it relies on the presence of quasi-rents in the exporter’s technology.
2.4 Pass-Through Characterization